OpenStudy (anonymous):

3 radical 5 x 2 radical 5?? simplify

OpenStudy (anonymous):

what did u get when u multiplied it

OpenStudy (anonymous):

multiply the outsides then the insides of the radicals

OpenStudy (anonymous):

6 radical 26

OpenStudy (anonymous):

25

OpenStudy (anonymous):

\[3\sqrt{5}\times 2\sqrt{5}=3\times 2\times \sqrt{5}^2=3\times 2\times 5\][

OpenStudy (anonymous):

Alight keep going @satellite73!!